# PDSS: A Privacy-Preserving Framework for Step-by-Step Distillation of Large Language Models

### v1

@@ Abstract
In the context of real-world applications, leveraging large language models (LLMs) for domain-specific tasks often faces two major challenges: domain-specific knowledge privacy and constrained resources. To address these issues, we propose PDSS, a privacy-preserving framework for step-by-step distillation of LLMs. PDSS works on a server-client architecture, wherein client transmits perturbed prompts to the server’s LLM for rationale generation. The generated rationales are then decoded by the client and used to enrich the training of task-specific small language model(SLM) within a multi-task learning paradigm. PDSS introduces two privacy protection strategies: the Exponential Mechanism Strategy and the Encoder-Decoder Strategy, balancing prompt privacy and rationale usability. Experiments demonstrate the effectiveness of PDSS in various text generation tasks, enabling the training of task-specific SLM with enhanced performance while prioritizing data privacy protection.
PDSS: A Privacy-Preserving Framework for Step-by-Step Distillation of Large Language Models
Tao Fan1, 2, Yan Kang2, Weijing Chen2, Hanlin Gu2, Yuanfeng Song2, Lixin Fan2, Kai Chen1, Qiang Yang1,2 1 Hong Kong University of Science and Technology, China 2 WeBank, China Correspondence: tfanac@cse.ust.hk
@@ 1 Introduction
Large Language Models(LLMs), boasting billions of parameters and remarkable text generation abilities, have risen as a revolutionary force in artificial intelligence. Prominent models, such as GPT-4 OpenAI (2023), LLaMATouvron et al. (2023), and QwenBai et al. (2023), have garnered the attention of researchers and practitioners alike, demonstrating unparalleled proficiency across numerous tasks. Nevertheless, the sheer size of these models presents significant obstacles for real-world deployment, particularly in environments with limited resources. Meanwhile, as LLMs gain escalating popularity and widespread utilization, privacy concerns have moved to the forefront, especially when it comes to user data and model inference. In contrast, Small Language Models(SLMs) often exhibit superior computational efficiency and faster convergence rates, rendering them perfectly suited for real-time applications or resource-constrained environments. Nonetheless, SLMs also possess certain drawbacks stemming from their performance limitations. The question then arises: How can we effectively combine the predictive prowess of LLMs with the nimbleness of SLMs, all while adhering to privacy requirements?
@@
To address these challenges, we introduce PDSS, a privacy-preserving framework for step-by-step distillation of LLMs. In our envisioned setup, there’s a high-powered server capable of deploying an LLM, paired with a client possessing more limited computational resources running SLM. The challenge lies in maintaining the privacy of client data while leveraging the server’s LLM to aid in training the client’s SLM for text generation tasks, thereby elevating its performance. PDSS aims to bridge this gap, enabling secure and efficient knowledge transfer between LLM and SLM, and ultimately enhancing the capabilities of the SLM without compromising privacy.
@@
As illustrated in Figure 1, within our framework, the process works as follows. Initially, the client transmits perturbed prompts to the server’s LLM, which are protected by the PDSS prompt encoder module, thus ensuring privacy protection. Subsequently, the server’s LLM generates perturbed rationales from these prompts through the Chain of Thought (COT) approach Wei et al. (2022) and relays them back to the client. Upon receiving these perturbed rationales, the client’s rationales decoder module reconstructs them into their original, aligned form corresponding to the raw prompt. Ultimately, the client incorporates these rationales as supplementary and enriching information for training its Task-Specific SLM within a multi-task learning paradigm Wei et al. (2022); Hsieh et al. (2023); Zhang and Yang (2021). These rationales justify the predicted labels and serve as insightful guidance for training smaller and domain-specific models.
@@
Within the PDSS framework, to achieve a balance between preserving the privacy of user prompts and enhancing the usability of rationales, we introduce two privacy protection strategies incorporated into the the prompt encoder module and the rationales decoder module: the Exponential Mechanism Strategy and the Encoder-Decoder Strategy. In the Exponential Mechanism Strategy, we utilize an exponential mechanism to obfuscate the prompts Tong et al. (2023), followed by decoding the perturbed rationales through In-Context Learning (ICL) Dong et al. (2022). In the Encoder-Decoder strategy, we utilize an Encoder-Decoder SLM specifically designed to encode raw prompts into perturbed prompts and subsequently decode perturbed rationales back into their original form. To effectively train this unified Encoder-Decoder SLM, we utilize a multi-task learning paradigm Zhang and Yang (2021), encompassing both the encoding and decoding training processes.
@@
  • Privacy-Preserving Framework for LLM Distillation. We propose PDSS, a novel framework that facilitates secure and efficient knowledge transfer from LLM to SLM in resource-constrained environments while adhering to privacy requirements. PDSS addresses the challenges posed by the massive size of LLMs for real-world deployment and the privacy concerns surrounding user data. By utilizing perturbed prompts and rationales, PDSS ensures data privacy while leveraging the predictive prowess of LLMs to enhance the performance of SLMs.
  • Innovative Privacy Protection Strategies. Within PDSS, we introduce two privacy protection strategies: the Exponential Mechanism Strategy and the Encoder-Decoder Strategy. The former utilizes an exponential mechanism to obfuscate user prompts, while the latter employs a specialized Encoder-Decoder SLM to encode and decode perturbed prompts and rationales. These strategies effectively balance user privacy and the usability of rationales, allowing for secure and enhanced training of the client’s SLM without compromising on privacy concerns.
  • Empirical Evaluation and Enhanced Performance of Task-Specific SLM. Through experiments on various text generation tasks, PDSS demonstrates the effectiveness of its framework in training task-specific SLM with enhanced performance. By harnessing the rationales generated by the server-side LLM, PDSS provides valuable task-specific knowledge to the SLM, enabling them to achieve significant improvements with the support of the LLM while prioritizing data privacy protections.
Figure 1: Overview of our proposed PDSS workflow.
Figure 2: Privacy-Preserving Rationals Generation Example.
@@ 2.1 Chain of Thought in Large Language Models
The Chain of Thought(COT) approach has recently garnered significant attention in the realm of LLMs, thanks primarily to its remarkable ability to enhance the reasoning capabilities of these models. This innovative concept was first introduced by Wei et al. (2022). Their research demonstrated that by prompting LLMs to produce a sequence of intermediary reasoning steps(rationales), the models’ performance in handling intricate reasoning tasks could be notably boosted. This groundbreaking study opened the door for further explorations into COT. Since the introduction of COT, several studies have delved into its extensions and variations. For example, Kojima et al. (2022) proposed the use of zero-shot COT, where the model is prompted to generate reasoning steps(rationales) without relying on prior examples. COT has also been applied to various domains, including arithmetic reasoningCobbe et al. (2021), commonsense reasoningKlein and Nabi (2020).
Nonetheless, despite the impressive feats achieved by LLMs, the adoption of LLMs in domain-specific applications with constrained resources poses a significant challengeFan et al. (2023) Kang et al. (2023). Recent studies by Hsieh et al. (2023) Ho et al. (2022) Li et al. (2023), have capitalized on the generated rationales as a form of insightful supervision to train smaller and domain-specific models. However, previous studies have not addressed the domain-specific data privacy issue that arises when LLMs and domain-specific smaller models are deployed across different parties. In our work, we endeavor to address this significant challenge.
2.2 Privacy Preserving LLM Inference
With the escalating popularity and widespread utilization of LLMs, privacy concerns have taken center stage, particularly regarding user data and model inference. Previous research efforts aimed at preserving privacy during LLM inference have predominantly focused on several key techniques, including differential privacy(DP) Dwork (2006), fully homomorphic encryption(FHE) Gentry (2009), and secure multi-party computation(MPC) Yao (1986) protocols.
Numerous studies have delved into the intricacies of LLM inference leveraging DP techniques. Notably, methods like SANTEXT+ Yue et al. (2021), CUSTEXT+ Chen et al. (2022), TextObfuscator Zhou et al. (2023) and InferDPT Tong et al. (2023) have harnessed differential privacy to sequentially replace sensitive words in the text with semantically similar alternatives from a predefined word adjacency list.
FHE and MPC techniques have also garnered attention as viable methods for ensuring privacy during LLM inference. For instance, CipherGPT Hou et al. (2023) proposes a secure matrix multiplication and a novel protocol for securely computing GELU within transformer architecture using FHE and MPC protocols to facilitate secure two-party GPT inference. Likewise, Puma Dong et al. (2023) has adopted FHE and MPC in its transformer architecture for secure third-party LLM inference. While FHE and MPC can be utilized for privacy-preserving text generation tasks, their practical applications remain limited primarily due to significant computational and communication overheads.
The advancements in privacy-preserving techniques, such as differential privacy, FHE, and MPC, offer promising solutions to mitigate privacy risks associated with LLM inference. However, balancing privacy and efficiency remains a challenge that requires further exploration and refinement.
3 The Proposed PDSS Framework
3.1 Overview
In this section, we introduce PDSS, an innovative privacy-preserving framework specifically designed for distilling step-by-step LLMs. The PDSS framework can enhance the performance of SLMs while maintaining privacy, leveraging the capabilities of LLM. We illustrate the PDSS in Figure 1 and describe the associated training algorithm in Algorithm 1. The workflow of PDSS is outlined as follows:
@@ 3.1 Overview
    In the client, Prompt Encoder Module perturbs these prompts before sending them to the server-side LLM.
@@
    In the server, the server-side LLM generates perturbed rationales based on these perturbed prompts and sends them back to the client.
  3. 3.
    In the client, Rationales Decoder Module decodes the perturbed rationales.
  4. 4.
    In the client, Task-Specific SLM Training Module employs both the original label data and the filter rationales data for multi-task learning.
3.2 Prompt Encoder Module
In the prompt encoder module, as illustrated in Figure 3, we propose two privacy protection strategies:
  1. 1.
    Exponential Mechanism Encoder Strategy. In the first strategy, we utilize an exponential mechanism McSherry and Talwar (2007)Tong et al. (2023), which satisfies the criteria for the $\epsilon-DP$. This strategy works by replacing each token in the prompt with a semantically similar one sampled from either a predetermined adjacency list or a randomly generated adjacency list, based on exponential mechanism.
    The Definition of Exponential Mechanism Tong et al. (2023). For a given scoring function $u:X\times Y\to R$, a randomized mechanism $M(X,u,Y)$ is $\epsilon-DP$ compliant if it satisfies:
    $P_{r}[y|x]\propto exp(\frac{\epsilon\cdot u(x,y)}{2\bigtriangleup u})$ (1)
    where the sensitivity $\bigtriangleup u$ is defined as:
    $\bigtriangleup u=\max_{x,x^{{}^{\prime}}\in X,y\in Y}|u(x,y)-u(x^{{}^{\prime}},y)|$ (2)
  2. 2.
    Encoder-Decoder Encoder Strategy. The tokens within a prompt differ significantly in terms of their importance and degree of privacy. Applying a uniform privacy budget $\epsilon$ across all tokens may not lead to the most optimal solution. To further optimize the privacy-utility balance, we propose an Encoder-Decoder strategy. This strategy is built upon the first exponential mechanism. In the Encoder-Decoder strategy, we utilize an Encoder-Decoder SLM specifically designed to encode raw prompts into perturbed prompts and subsequently decode perturbed rationales back into their original form. This strategy involves two training process: encoding training process and decoding training process. In this section, we mainly focus on encoding training process, as illustrated in Figure 3.
@@ 3.2 Prompt Encoder Module
    The Encoder objective can be formulated as follows:
@@
As illustrated in Figure 2, we can observe an exemplary comparison between the original input and its perturbed input in Step 1 and Step 2. This perturbed prompt serves as the new, privacy-enhanced input for further processing.
By incorporating this perturbation mechanism, we ensure that the privacy of the original prompt is preserved. This approach not only satisfies the privacy requirements but also enables effective data utilization for downstream tasks, striking a balance between privacy and utility.
Figure 3: Prompt Encoder Module.
@@ 3.3 Generating Perturbed Rationales from LLM
When the server-side LLM receives the perturbed prompt, we leverage the Chain-of-Thought (CoT) prompting technique introduced by Wei et al. (2022) to generate rationales from the LLM using this perturbed prompt. These generated rationales, which are also perturbed, are then transmitted to the client. For instance, as illustrated in Figure 2, given a perturbed prompt in the Step 2, the LLM generates perturbed rationales in the Step 3.
3.4 Rationales Decoder Module
Once the client receives the perturbed rationales from the server-side LLM, it must initiate a "decoder" process within the rationales decoder module to decode the rationales. In rationales decoder module, as illustrated in Figure 4, we also propose two strategies correspond to the two protection strategy of the prompt encoder module:

### v2

# FedCoT: Federated Chain-of-Thought Distillation for Large Language Models

@@ Abstract
Large Language Models (LLMs) have emerged as a transformative force in artificial intelligence, demonstrating exceptional proficiency across various tasks. However, their deployment in resource-constrained environments and concerns over user data privacy pose significant challenges. In contrast, Small Language Models (SLMs) offer computational efficiency but often lag in performance. To address these issues, we propose FedCoT, a federated framework designed for the Chain-of-Thought (CoT) distillation of knowledge from LLMs to SLMs, while ensuring the preservation of clients’ data privacy. FedCoT ensures secure and efficient knowledge transfer from an LLM on a high-powered server to an SLM on a resource-constrained client, while adhering to privacy requirements. Leveraging perturbed prompts and rationales generated through the CoT approach, the framework enhances the performance of the client’s SLM without compromising user data privacy within a multi-task learning framework. We propose two privacy protection strategies: the Exponential Mechanism Strategy and the Adaptive Exponential Mechanism Strategy, which balance user prompt privacy and the usability of rationales. Empirical evaluation on various text generation tasks demonstrates the effectiveness of FedCoT in training task-specific SLMs with enhanced performance while prioritizing data privacy protection. Our code has been contributed to the FATE open-source project and is now publicly accessible at https://github.com/FederatedAI/FATE-LLM/tree/main/python/fate_llm/algo/fedcot
FedCoT: Federated Chain-of-Thought Distillation for Large Language Models
Tao Fan1, 2, Weijing Chen2, Yan Kang2, Guoqiang Ma2, Hanlin Gu2, Yuanfeng Song2, Lixin Fan2, Qiang Yang3 1 Hong Kong University of Science and Technology, Hong Kong, China 2 WeBank, China 3Hong Kong Polytechnic University, Hong Kong, China Correspondence: tfanac@cse.ust.hk, qyang@cse.ust.hk
@@ 1 Introduction
Large Language Models (LLMs) have risen as a revolutionary force in artificial intelligence. Prominent LLMs, such as GPT-4 Gpt-4, LLaMA touvron2023llama, and Qwen bai2023qwen, have garnered the attention of researchers and practitioners alike, demonstrating unparalleled proficiency across numerous tasks. Nevertheless, the sheer size of these models presents significant obstacles for real-world deployment, particularly in environments with limited resources fan2025ten; fan2025fedmkt; fan2023fate-llm; kang2023grounding. Meanwhile, as LLMs gain escalating popularity and widespread utilization, privacy concerns have moved to the forefront, especially when it comes to user data and LLMs inference. In contrast, Small Language Models (SLMs) often exhibit superior computational efficiency and faster convergence rates, rendering them perfectly suited for real-time applications or resource-constrained environments. Nonetheless, SLMs also possess certain drawbacks stemming from their performance limitations. The question then arises: How can we effectively combine the predictive prowess of LLMs with the nimbleness of SLMs, all while adhering to privacy requirements?
@@
To address these challenges, we propose FedCoT, a federated framework designed for the Chain-of-Thought (CoT) wei2022chain distillation of knowledge from LLMs to SLMs, while ensuring the preservation of clients’ data privacy. FedCoT ensures secure and efficient knowledge transfer from an LLM on a high-powered server to an SLM on a resource-constrained client. The challenge lies in maintaining the privacy of client data while leveraging the server’s LLM to aid in training the client’s SLM for text generation tasks, thereby elevating its performance. FedCoT aims to bridge this gap, enabling secure and efficient knowledge transfer between LLM and SLM, and ultimately enhancing the capabilities of the SLM without compromising privacy.
@@
As illustrated in Figure 1(a), within our framework, the process works as follows. Initially, the client transmits perturbed prompts to the server’s LLM. These prompts are protected by the FedCoT prompt encoder, which employs Differential Privacy (DP) principles dwork2006differential; mcsherry2007mechanism, ensuring privacy protection. Subsequently, the server’s LLM generates perturbed rationales from these prompts through the CoT approach and relays them back to the client. Upon receiving these perturbed rationales, the client’s rationales decoder reconstructs them into their original, aligned form corresponding to the raw prompt. Ultimately, the client utilizes CoT knowledge distillation hsieh2023distilling; li2023symbolic to train its Task-Specific SLM. This process leverages both label data and rationales within a multi-task learning paradigm wei2022chain; hsieh2023distilling; zhang2021survey. These rationales justify the predicted labels and serve as insightful guidance for training smaller and domain-specific models.
Previous endeavors to incorporate DP into language models, specifically through DP-SGD song2013stochastic, have primarily centered on navigating the delicate balance between utility and privacy. This is achieved by introducing calibrated noise into gradients or text representations during the model training process. Nonetheless, these methods inherently rely on a trusted server to gather data from data owners for model training chen-etal-2023-customized, significantly limiting their applicability in scenarios where such trusted servers are not available, as is the case in our research context.
@@
Within the FedCoT framework, to achieve a balance between preserving the privacy of user prompts and enhancing the usability of rationales, we introduce two privacy protection strategies: the Exponential Mechanism Strategy and the Adaptive Exponential Mechanism Strategy. In the Exponential Mechanism Strategy, we utilize an exponential mechanism to obfuscate the prompts mcsherry2007mechanism; yue2021differential; chen-etal-2023-customized, followed by decoding the perturbed rationales through In-Context Learning (ICL) dong2024survey; tong2025inferdpt. In the Adaptive Exponential Mechanism Strategy, we utilize an Encoder-Decoder SLM specifically designed to encode original prompts into perturbed prompts and subsequently decode perturbed rationales back into their original form. To effectively train this unified Encoder-Decoder SLM, we utilize a multi-task learning paradigm zhang2021survey, encompassing both the encoding and decoding training processes.
@@
  • Federated Framework for CoT Distillation in LLMs. We propose FedCoT, a novel federated framework that facilitates secure and efficient knowledge transfer from LLM to SLM in resource-constrained environments. FedCoT leverages CoT knowledge distillation to enhance Task-Specific SLM within the client. This process leverages rationales produced by the LLM on the server, thereby enriching the client-side SLMs with valuable task-related knowledge.
  • Privacy as a Priority. FedCoT leverages an Adaptive Exponential Mechanism Strategy tailored for encoding prompt to ensure their obfuscation and decoding perturbed rationales. The strategies effectively balance user prompt privacy and the usability of rationales.
  • Empirical Evaluation and Enhanced Performance of Task-Specific SLM. Through experiments on various text generation tasks, FedCoT demonstrates the effectiveness of its framework in training task-specific SLM with enhanced performance. By harnessing the rationales generated by the server-side LLM, FedCoT provides valuable task-specific knowledge to the SLM.
(a) Overview of our proposed FedCoT framework.
(b) Privacy-Preserving Rationals Generation.
Figure 1: The overview of our proposed FedCoT. The FedCoT comprises four key components: (1) The Prompt Encoder, which perturbs user prompts to ensure privacy; (2) The LLM, generating perturbed rationales based on the perturbed prompts; (3) The Perturbed Rationales Decoder, which decodes the perturbed rationales back into a usable form; (4) The Task-Specific SLM Enhancing via CoT Knowledge Distillation, utilizing both original labeled data and filtered rationales data for multi-task learning.
@@ 2 Related Work
2.1 Differential Privacy
In this section, We briefly revisit two important definitions of differential privacy: $\epsilon$-Differential Privacy and Exponential Mechanism (EM).
$\epsilon$-Differential Privacy. Differential privacy (DP)  dwork2006differential is a rigorous mathematical framework that provides strong privacy guarantees for data analysis. It ensures that the output of an algorithm remains statistically indistinguishable whether a particular individual’s data is included or excluded from the dataset. Formally, a randomized mechanism $M$ provides $\epsilon$-differential privacy if for all neighboring datasets $D$ and $D^{\prime}$ (differing in at most one record) and for all sets $S$ of possible outputs:
$\Pr[M(D)\in S]\leq e^{\epsilon}\Pr[M(D^{\prime})\in S]$ (1)
where $\epsilon$ is the privacy budget that controls the level of indistinguishability.
Exponential Mechanism. The Exponential Mechanism  mcsherry2007mechanism allows for the selection of an outcome from a set of possible outcomes with probabilities proportional to the exponential of their utility scores. Formally, given a utility function $u:D\times R\rightarrow\mathbb{R}$ that maps each dataset $D$ and possible outcome $r$ to a real-valued score, the Exponential Mechanism $M(D,u,R)$ satisfies $\epsilon$-differential privacy if it selects and outputs an $r\in R$ with probability:
$\Pr[M(D)=r]\propto\exp\left(\frac{\epsilon u(D,r)}{2\Delta u}\right)$ (2)
where $\Delta u$ is the sensitivity of the utility function (in our work, we use cosine similarity as the utility function), defined as the maximum change in utility score when a single record is added or removed from the dataset:
$\Delta u=\max_{D,D^{\prime},r}|u(D,r)-u(D^{\prime},r)|$ (3)
@@ 2.2 Chain of Thought in Large Language Models
The Chain of Thought (CoT) approach has recently garnered significant attention in the realm of LLMs, thanks primarily to its remarkable ability to enhance the reasoning capabilities of these models. This innovative concept was first introduced by  wei2022chain. Their research demonstrated that by prompting LLMs to produce a sequence of intermediary reasoning steps (rationales), the models’ performance in handling intricate reasoning tasks could be notably boosted. Since the introduction of CoT, several studies have delved into its extensions and variations. For example,  kojima2022large proposed the use of zero-shot CoT, where the model is prompted to generate rationales without relying on prior examples. CoT has also been applied to various domains, including arithmetic reasoning cobbe2021training, commonsense reasoning klein2020contrastive. Recent studies by hsieh2023distilling; ho2023large; li2023symbolic, have capitalized on the generated rationales as a form of insightful supervision to train smaller and domain-specific models. However, previous studies have not addressed the domain-specific data privacy issue that arises when LLMs and domain-specific smaller models are deployed across different parties. In our work, we endeavor to address this significant challenge.
3 The Proposed FedCoT Framework
In this section, we introduce FedCoT, a federated framework designed for the CoT distillation of knowledge from LLMs hosted on a high-powered server to SLMs deployed on a resource-constrained client. The FedCoT framework can enhance the performance of SLMs while maintaining client data’s privacy, leveraging the capabilities of LLM. We assume the server to be semi-honest, implying that it may attempt to recover the private data of the client from the information it observes. We illustrate the FedCoT in Figure 1(a), outline its training algorithm in Algorithm 1, and detail its resource requirements in Appendix A.
Algorithm 1 FedCoT
Input:
$T$: total number of rounds;
$\mathcal{P}$: encoding training datasets;
$\mathcal{R}$: decoding training datasets;
$\mathcal{D}$: task-specific training datasets;
$\eta_{\phi}$: learning rate of Encoder-Decoder SLM;
$\eta_{\omega}$: learning rate of Task-Specific SLM.
Output: $g_{\phi}$, $f_{\omega}$.
[…]
2:for each epoch $t\in[T]$ do
3:  $\phi^{t+1}\leftarrow\phi^{t}-\eta_{\phi}\nabla\mathcal{L}_{1}$.
4:end for
[…]
12:for each epoch $t\in[T]$ do
13:  $\omega^{t+1}\leftarrow\omega^{t}-\eta_{\omega}\nabla\mathcal{L}_{2}$.
14:end for
3.1 Privacy Preserving Prompt Encoder
Before the client transmits its raw prompts to the server-side LLM, we need the privacy protection strategy to protect the raw prompts. In this section, we propose two privacy protection strategies:
@@ 3.1 Privacy Preserving Prompt Encoder
    Exponential Mechanism Encoder Strategy. In the first strategy, we utilize an exponential mechanism mcsherry2007mechanism, which satisfies the criteria for the $\epsilon$-DP. For detailed information about the exponential mechanism, please refer to Section 2.1.
    Let us consider an Exponential Mechanism $M(\cdot)$. Given a input prompt $p=\{x_{i}\}_{i=1}^{S}$ comprising $S$ tokens, a set $X$ encompassing all possible input tokens, and a set $Y$ of all potential output tokens, the mechanism $M(\cdot)$ is applied to each input token $x_{i}\in p$. If $x_{i}$ belongs to $X$, it is replaced with an output token $y_{i}$ from $Y$. Through this process, we obtain a perturbed prompt $p^{p}=\{y_{i}\}_{i=1}^{S}$.
@@
    Adaptive Exponential Mechanism Encoder Strategy. The tokens within a prompt differ significantly in terms of their importance and degree of privacy. Applying a uniform privacy budget $\epsilon$ across all tokens may not lead to the most optimal solution. To further optimize the privacy-utility balance, we propose an Adaptive Exponential Mechanism Encoder strategy. This strategy is built upon the first exponential mechanism. In the Adaptive Exponential Mechanism Encoder strategy, we utilize an Encoder-Decoder SLM specifically designed to encode raw prompts into perturbed prompts and subsequently decode perturbed rationales back into their original form. This strategy involves two training process: encoding training process and decoding training process. In this section, we mainly focus on encoding training process.
@@
    The Prompt Encoder objective can be formulated as follows:
@@
As illustrated in Figure1(b), we can observe an exemplary comparison between the original prompt and its perturbed prompt in Step 1 and Step 2. This perturbed prompt serves as the new, privacy-enhanced input for further processing.
@@ 3.2 Generating Perturbed Rationales from LLM
When the server-side LLM receives the perturbed prompt, we leverage the Chain-of-Thought (CoT) prompting technique introduced by wei2022chain to generate rationales from the LLM using this perturbed prompt. These generated rationales, which are also perturbed, are then transmitted to the client. For instance, as illustrated in Figure 1(b), given a perturbed prompt in the Step 2, the LLM generates perturbed rationales in the Step 3.
3.3 Perturbed Rationales Decoder
Once the client receives the perturbed rationales from the server-side LLM, it must initiate a "decoder" process to decode the rationales. In this section, we also propose two strategies correspond to the two protection strategy of the prompt encoder module: